# TimeDRL: Disentangled Representation Learning for Multivariate Time-Series

### 07 Dec 2023 v1

@@ V-D2 Data Augmentations
The foundational principle of TimeDRL is to avoid any data augmentation to prevent the introduction of inductive biases. Accordingly, our design of both the timestamp-predictive task and the instance-contrastive task does not involve any augmentation methods. In our experiment, we aim to demonstrate the potential drawbacks of ignoring the issue of inductive bias. In Table VII, we experiment with 6 time-series-specific data augmentation methods [52, 12]. Jittering simulates sensor noise through additive Gaussian noise. Scaling adjusts the magnitude of data by multiplying it with a random scalar. Rotation modifies the dataset by permuting the order of features and potentially flipping the sign of feature values. Permutation slices the data into segments, then randomly permutes these segments to create a new time-series instance. Masking randomly sets values in the time-series data to zero. Cropping removes the left and right regions of a time-series instance and fills the gaps with zeros to maintain the same sequence length.
In the experiment detailed in Table VII, the application of any augmentation method led to a decline in performance, with an average 27.77% decrease in MSE. The most significant drop was observed with Rotation augmentation, where MSE decreased by 68.15% and MAE decreased by 33.2%. TS2Vec [12] addresses the inductive bias issue but still employs Masking and Cropping augmentations. Our results indicate that these two methods are relatively less harmful compared to others, yet they still lead to performance degradation. This experiment supports our initial assumption that the complete avoidance of augmentation methods is crucial for eliminating inductive bias, thereby ensuring the optimal performance of TimeDRL.
Table VII: Ablation study on data augmentation. We use the prediction length $T=168$ for ETTh1. The best results are in bold.
@@
0.462 (-10.57%)
0.453 (-6.09%)
Table VIII: Ablation study on pooling methods. The best results are in bold.
Pooling
Method
@@ V-D3 Pooling Methods
In TimeDRL, we employ a dedicated [CLS] token strategy to derive instance-level embeddings directly from patched timestamp-level data. Despite this approach, we recognize the theoretical possibility of extracting instance-level embeddings from timestamp-level embeddings using pooling methods. To explore this, we conducted experiments with 3 other different pooling strategies for instance-level embeddings, as detailed in Table VII. Last utilizes the last timestamp-level embedding as the instance-level representation. GAP employs global average pooling, averaging the timestamp-level embeddings across the time axis to aggregate the instance-level embedding. All flattens all timestamp-level embeddings to create a singular instance-level representation.
The results in Table VIII reveal that the [CLS] token strategy employed by TimeDRL outperforms other methods, achieving an average increase of 13.02% in accuracy. The least effective pooling method is GAP, commonly used in the time-series domain [12]. This method experiences the most significant performance drop due to the anisotropy problem, with accuracy decreasing by 19.05%. These results demonstrate the importance of disentangling timestamp-level and instance-level embeddings, a key factor contributing to TimeDRL’s superior performance compared to other baseline methods.
Table IX: Ablation study on the architecture of the backbone encoder. We use the prediction length $T=168$ for ETTh1. The best results are in bold.
@@
MSE
MAE
Transformer Encoder (Ours) 0.418 0.427
@@
0.443 (-5.92%)
0.443 (-3.69%)
Table X: Ablation study on the stop gradient operation. The best results are in bold.
Stop
Gradient
@@ V-D4 Encoder Architectures
Transformers are well-known for their success in downstream time-series tasks [35, 36], but for self-supervised learning in time-series, CNN-based [24, 12] and RNN-based [14] models are usually chosen over Transformers. TimeDRL is designed to make the best use of Transformers in self-supervised learning for time-series data, aiming to show the Transformer’s strong capabilities in this field. Within TimeDRL, the Transformer Encoder is utilized as the core encoder. To benchmark its performance against other encoder architectures, we conducted experiments with 5 different models, detailed in Table IX. Transformer Decoder employs a similar architecture to the Transformer Encoder, with the key distinction of using masked self-attention. This ensures that each timestamp’s embedding attends only to preceding timestamps, not subsequent ones. ResNet adopts the well-known architecture from computer vision’s ResNet18, modifying it with one-dimensional convolutions suitable for time-series data. TCN [49] combines dilations and residual connections with causal convolutions, specifically tailored for autoregressive prediction in time-series. LSTM uses Long Short-Term Memory units for capturing dependencies in sequential data. A uni-directional LSTM is utilized, focusing on past and present data to prevent future data leakage. Bi-LSTM follows the LSTM structure but incorporates bi-directional processing, allowing the model to integrate information from both past and future timestamps.
The findings presented in Table IX indicate that the Transformer Encoder, when combined with our two pretext tasks, is the most effective model architecture. It demonstrates superior performance, with an average increase of 13.91% in MSE and 8.38% in MAE over other architectures. In contrast, the Transformer Decoder shows a decrease in performance with MSE reduced by 11.26%. This difference highlights the critical role of bidirectional self-attention in achieving a comprehensive understanding of the entire sequence. Similarly, when comparing LSTM with Bi-LSTM, the latter exhibits enhanced performance due to its ability to access both past and future information. These findings highlight the importance of having full temporal access for each timestamp and confirm the Transformer Encoder’s strong capability of acquiring effective time-series representations.
@@ V-D5 Stop Gradient
To address sampling bias, our approach includes an extra prediction head coupled with a stop-gradient operation. This asymmetric design, featuring one pathway with the additional prediction head and another with the stop-gradient, proves effective in preventing model collapse, as demonstrated in studies like [11, 42]. The results in Table X show that eliminating the stop-gradient operation leads to a significant 11.11% drop in accuracy, underscoring the crucial function of the stop-gradient element in this asymmetric architecture.
@@ VI Conclusion
This paper introduces TimeDRL, a novel multivariate time-series representation learning framework with disentangled dual-level embeddings. Our framework disentangles timestamp-level and instance-level embeddings, making it versatile for various time-series tasks such as forecasting and classification. At its core, TimeDRL utilizes a [CLS] token strategy to extract contextualized instance-level embeddings from patched timestamp-level data. Two pretext tasks are introduced to optimize representations: the timestamp-predictive task focuses on optimizing timestamp-level embeddings with predictive loss, while the instance-contrastive task aims to optimize instance-level embeddings with contrastive loss. To address the challenge of inductive bias, TimeDRL avoids the direct application of any data augmentations in both pretext tasks. In the timestamp-predictive task, TimeDRL relies on the reconstruction error of patched time-series data, deliberately avoiding any input masking. For the instance-contrastive task, TimeDRL exploits the randomness of dropout layers to create two distinct embedding views from a single data input. Comprehensive experiments on 6 forecasting datasets and 5 classification datasets demonstrate TimeDRL’s superior performance, delivering an average improvement of 57.98% in MSE for forecasting and 1.25% in accuracy for classification. TimeDRL also supports the effective capability with limited labeled data in semi-supervised learning scenarios. Moreover, extensive ablation studies with detailed analysis further validated the relative contributions of each component within TimeDRL’s architecture.

### 13 Mar 2024 v2

@@ V-D2 Data Augmentations
The foundational principle of TimeDRL is to avoid any data augmentation to prevent the introduction of inductive biases. Accordingly, our design of both the timestamp-predictive task and the instance-contrastive task does not involve any augmentation methods. In our experiment, we aim to demonstrate the potential drawbacks of ignoring the issue of inductive bias. In Table VI, we experiment with 6 time-series-specific data augmentation methods [52, 12]. Jittering simulates sensor noise through additive Gaussian noise. Scaling adjusts the magnitude of data by multiplying it with a random scalar. Rotation modifies the dataset by permuting the order of features and potentially flipping the sign of feature values. Permutation slices the data into segments, then randomly permutes these segments to create a new time-series instance. Masking randomly sets values in the time-series data to zero. Cropping removes the left and right regions of a time-series instance and fills the gaps with zeros to maintain the same sequence length.
In the experiment detailed in Table VI, the application of any augmentation method led to a decline in performance, with an average increase in MSE of 27.77% for the ETTh1 dataset and 57.37% for the Exchange dataset. The most significant deterioration was observed with Rotation augmentation, where MSE increased by 68.15% for the ETTh1 dataset and 174.46% for the Exchange dataset. TS2Vec [12] addresses the inductive bias issue but still employs Masking and Cropping augmentations. Our results indicate that these two methods are relatively less harmful compared to others, yet they still lead to performance degradation. This experiment supports our initial assumption that the complete avoidance of augmentation methods is crucial for eliminating inductive bias, thereby ensuring the optimal performance of TimeDRL.
Table VI: Ablation study on data augmentation. We use the prediction length $T=168$. The best results are in bold.
@@
0.462 (+10.57%)
0.216 (+47.70%)
Table VII: Ablation study on pooling methods. The best results are in bold.
Pooling Method
@@ V-D3 Pooling Methods
In TimeDRL, we employ a dedicated [CLS] token strategy to derive instance-level embeddings directly from patched time-series data. Despite this approach, we recognize the theoretical possibility of extracting instance-level embeddings from timestamp-level embeddings using pooling methods. To explore this, we conducted experiments with 3 other different pooling strategies for instance-level embeddings, as detailed in Table VI. Last utilizes the last timestamp-level embedding as the instance-level representation. GAP employs global average pooling, averaging the timestamp-level embeddings across the time axis to aggregate the instance-level embedding. All flattens all timestamp-level embeddings to create a singular instance-level representation.
The results in Table VII reveal that employing strategies other than the [CLS] token strategy used by TimeDRL results in an average decrease in accuracy of 11.11% for the FingerMovements dataset and 16.75% for the Epilepsy dataset, highlighting the superior performance of the [CLS] token approach. The least effective pooling method is GAP, commonly used in the time-series domain [12]. This method experiences the most significant performance drop due to the anisotropy problem. These results demonstrate the importance of disentangling timestamp-level and instance-level embeddings, a key factor contributing to TimeDRL’s superior performance compared to other baseline methods.
Table VIII: Ablation study on the architecture of the backbone encoder. We use the prediction length $T=168$. The best results are in bold.
@@
Exchange
Transformer Encoder (Ours) 0.418 0.146
@@
0.443 (+5.92%)
0.153 (+4.62%)
Table IX: Ablation study on the stop gradient operation. The best results are in bold.
Stop Gradient
@@ V-D4 Encoder Architectures
Transformers are well-known for their success in downstream time-series tasks [35, 36], but for self-supervised learning in time-series, CNN-based [24, 12] and RNN-based [14] models are usually chosen over Transformers. TimeDRL is designed to make the best use of Transformers in self-supervised learning for time-series data, aiming to show the Transformer’s strong capabilities in this field. Within TimeDRL, the Transformer encoder is utilized as the core encoder. To benchmark its performance against other encoder architectures, we conducted experiments with 5 different models, detailed in Table VIII. Transformer Decoder employs a similar architecture to the Transformer encoder, with the key distinction of using masked self-attention. This ensures that each timestamp’s embedding attends only to preceding timestamps, not subsequent ones. ResNet adopts the well-known architecture from computer vision’s ResNet18, modifying it with one-dimensional convolutions suitable for time-series data. TCN [49] combines dilations and residual connections with causal convolutions, specifically tailored for autoregressive prediction in time-series. LSTM uses Long Short-Term Memory units for capturing dependencies in sequential data. A uni-directional LSTM is utilized, focusing on past and present data to prevent future data leakage. Bi-LSTM follows the LSTM structure but incorporates bi-directional processing, allowing the model to integrate information from both past and future timestamps.
The findings presented in Table VIII indicate that not using the Transformer encoder, in combination with our two pretext tasks, results in worse model performance. Specifically, this results in an average increase of 17.30% in MSE for the ETTh1 dataset and 6.60% for the Exchange dataset compared to when the Transformer encoder is utilized, underscoring its efficacy over other architectures. In contrast, compared to the Transformer encoder, the Transformer Decoder shows a decline in performance, with an 11.26% increase in MSE for the ETTh1 dataset and 8.28% for the Exchange dataset. This difference highlights the critical role of bidirectional self-attention in achieving a comprehensive understanding of the entire sequence. Similarly, when comparing LSTM with Bi-LSTM, the latter exhibits enhanced performance due to its ability to access both past and future information. These findings highlight the importance of having full temporal access for each timestamp and confirm the Transformer encoder’s strong capability of acquiring effective time-series representations.
@@ V-D5 Stop Gradient
To address sampling bias, our approach includes an extra prediction head coupled with a stop-gradient operation. This asymmetric design, featuring one pathway with the additional prediction head and another with the stop-gradient, proves effective in preventing model collapse, as demonstrated in studies like [11, 42]. The results in Table IX show that eliminating the stop-gradient operation leads to a significant drop in accuracy of 11.11% for the FingerMovements dataset and 16.75% for the Epilepsy dataset, underscoring the crucial function of the stop-gradient element in this asymmetric architecture.
@@ VI Conclusion
This paper introduces TimeDRL, a novel multivariate time-series representation learning framework with disentangled dual-level embeddings. Our framework disentangles timestamp-level and instance-level embeddings, making it versatile for various time-series tasks such as forecasting and classification. At its core, TimeDRL utilizes a [CLS] token strategy to extract contextualized instance-level embeddings from patched time-series data. Two pretext tasks are introduced to optimize representations: the timestamp-predictive task focuses on optimizing timestamp-level embeddings with predictive loss, while the instance-contrastive task aims to optimize instance-level embeddings with contrastive loss. To address the challenge of inductive bias, TimeDRL avoids the direct application of any data augmentations in both pretext tasks. In the timestamp-predictive task, TimeDRL relies on the reconstruction error of patched time-series data, deliberately avoiding any input masking. For the instance-contrastive task, TimeDRL exploits the randomness of dropout layers to create two distinct embedding views from a single data input. Comprehensive experiments on 6 forecasting datasets and 5 classification datasets demonstrate TimeDRL’s superior performance, delivering an average improvement of 58.02% in MSE for forecasting and 1.48% in accuracy for classification. TimeDRL also supports the effective capability with limited labeled data in semi-supervised learning scenarios. Moreover, extensive ablation studies with detailed analysis further validated the relative contributions of each component within TimeDRL’s architecture. In future work, we will enhance TimeDRL for time-series classification toward a more comprehensive foundation model. Moreover, the comparisons with large language model (LLM)-based approaches will be investigated to study the impacts of LLMs in the time-series domain.